# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Fundamental Limits

### v1

@@ IV Numerical Analysis
Figure 1: The causal tradeoff curves when $\lambda=0.1$ and $\rho\in\{0.0,0.3,0.7,1.0\}$.
@@
Figure 2: The causal tradeoff curves when $\lambda=0.3$ and $\rho\in\{0.0,0.3,0.7,1.0\}$.
V Conclusions
In this article, we established a theoretical framework for networked estimation over unreliable channels with unreliable feedback. We characterized a globally optimal coding policy profile consisting of a symmetric threshold scheduling policy for the encoder and a linear estimation policy for the decoder. In addition, we obtained the recursive equations that need to be solved by the encoder and the decoder, and specified the computational complexity required for the implementation of the optimal policies. Finally, we provided a numerical analysis to demonstrate and compare the fundamental performance limits of the networked system in different communication regimes.
@@ V Conclusions
This section is dedicated to the derivation of the main results. The following Lemma shows that the conditional means $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ and $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$ are instrumental for the operations performed by the encoder and the decoder.
@@
Let $\epsilon^{o}$ be a globally optimal scheduling policy that is implemented. Given all information available at the decoder, $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$ is the optimal value that minimizes the mean square error at the decoder at time $k$. Moreover, $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ fuses all the previous and current outputs of the process that are available at the encoder at time $k$. This implies that from the minimum mean-square-error perspective the decoder is able to develop a state estimate upon the receipt of a message at time $k+1$, that would be the same if it had all the previous outputs of the process until time $k$, which is the best possible case for the decoder. Hence, both claims hold. ∎
The next two lemmas characterize the optimal estimators at the encoder and the decoder.
@@
The optimal estimator minimizing the mean square error at the encoder satisfies the recursive equations
$\check{x}_{k}$ $=m_{k}+K_{k}(y_{k}-C_{k}m_{k})$ (20)
$m_{k}$ $=A_{k-1}\check{x}_{k-1}$ (21)
$Q_{k}$ $=(M_{k}^{-1}+C_{k}^{T}V_{k}^{-1}C_{k})^{-1}$ (22)
$M_{k}$ $=A_{k-1}Q_{k-1}A_{k-1}^{T}+W_{k-1}$ (23)
@@
The optimal estimator minimizing the mean square error at the decoder satisfies the recursive equation
$\hat{x}_{k}$ $=A_{k-1}\hat{x}_{k-1}+\delta_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1})$
$\quad+(1-\delta_{k-1})\imath_{k-1}+\delta_{k-1}(1-\gamma_{k-1})\jmath_{k-1}$ (24)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}$, where $\imath_{k-1}=A_{k-1}\operatorname{\mathsf{E}}[\hat{e}_{k-1}|\mathcal{I}_{k-1}^{d},\delta_{k-1}=0]$ and $\jmath_{k-1}=A_{k-1}\operatorname{\mathsf{E}}[\hat{e}_{k-1}|\mathcal{I}_{k-1}^{d},\delta_{k-1}=1]$.
@@
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\operatorname{\mathsf{E}}\Big{[}x_{k-1}\big{|}\mathcal{I}_{k}^{d}\Big{]}$ (25)
for $k\in\mathbb{N}_{[1,N]}$ as $w_{k-1}$ is independent of $\mathcal{I}^{d}_{k}$ and has zero mean. When sensory information is successfully delivered in the forward channel at time $k$, we have $z^{f}_{k}=\check{x}_{k-1}$. In this case, we get $\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k}]=\operatorname{\mathsf{E}}[x_{k-1}|\check{x}_{k-1},Q_{k-1}]=\check{x}_{k-1}$ as $\{\check{x}_{k-1},Q_{k-1}\}$ is a sufficient statistic of $\mathcal{I}^{e}_{k}$. Hence, using (25), when $z^{f}_{k}=\check{x}_{k-1}$, i.e., when $\delta_{k-1}=1\wedge\gamma_{k-1}=1$, we get
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\check{x}_{k-1}$ (26)
for $k\in\mathbb{N}_{[1,N]}$. However, when no sensory information is delivered in the forward channel at time $k$, we have $z^{f}_{k}=\mathfrak{E}$ or $z^{f}_{k}=\mathfrak{F}$. Define $p_{k-1}:=\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1},\delta_{k-1}=0]-\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1}]$ when $z^{f}_{k}=\mathfrak{E}$, and $q_{k-1}:=\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1},\delta_{k-1}=1]-\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1}]$ when $z^{f}_{k}=\mathfrak{F}$. Then, using (25) and the definition of $p_{k-1}$, when $z^{f}_{k}=\mathfrak{E}$, i.e., when $\delta_{k-1}=0$, we get
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\operatorname{\mathsf{E}}\Big{[}x_{k-1}\big{|}\mathcal{I}^{d}_{k-1}\Big{]}+A_{k-1}p_{k-1}$ (27)
for $k\in\mathbb{N}_{[1,N]}$, where we used the fact that $\{\mathcal{I}^{d}_{k-1},z^{f}_{k}=\mathfrak{E}\}$ is a sufficient statistic of $\mathcal{I}^{d}_{k}$. Moreover, using (25) and the definition of $q_{k-1}$, when $z^{f}_{k}=\mathfrak{F}$, i.e., when $\delta_{k-1}=1\wedge\gamma_{k-1}=0$, we get
$\operatorname{\mathsf{E}}\Big{[}x_{k}\big{|}\mathcal{I}^{d}_{k}\Big{]}=A_{k-1}\operatorname{\mathsf{E}}\Big{[}x_{k-1}\big{|}\mathcal{I}^{d}_{k-1}\Big{]}+A_{k-1}q_{k-1}$ (28)
for $k\in\mathbb{N}_{[1,N]}$, where we used the facts that $\{\mathcal{I}^{d}_{k-1},z^{f}_{k}=\mathfrak{G}\}$ is a sufficient statistic of $\mathcal{I}^{d}_{k}$ and that $x_{k-1}$ is independent of $\gamma_{k-1}$. Now, define $\imath_{k-1}:=A_{k-1}p_{k-1}$ and $\jmath_{k-1}:=A_{k-1}q_{k-1}$ as the signaling residuals. We can obtain (3) by combining (26), (27), and (28). Finally, the initial condition is $\operatorname{\mathsf{E}}[x_{0}]=m_{0}$, as no information is available at time $k=0$. ∎
The next two lemmas provide certain properties pertaining to the estimation mismatch and the signaling residuals $\imath_{k}$ and $\jmath_{k}$.
@@
$\tilde{e}_{k}$ $=(1-\delta_{k-1}\gamma_{k-1})A_{k-1}\tilde{e}_{k-1}+K_{k}\nu_{k}+\mu_{k-1}$ (29)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\tilde{e}_{0}=K_{0}\nu_{0}$, where $\mu_{k-1}=-(1-\delta_{k-1})\imath_{k-1}-\delta_{k-1}(1-\gamma_{k-1})\jmath_{k-1}$. Besides, the conditional mean and the conditional covariance of the estimation mismatch at the encoder satisfy the recursive equations
$\breve{e}_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}\breve{e}_{k-1}+K_{k}\nu_{k}+\varsigma_{k-1}$ (30)
@@
$\quad+\bar{\bar{\theta}}_{k-1}A_{k-1}\breve{e}_{k-1}\breve{e}_{k-1}^{T}A_{k-1}^{T}+\Xi_{k-1}$ (31)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\breve{e}_{0}=K_{0}\nu_{0}$ and $R_{0}=0$, where $\bar{\theta}_{k-1}=\theta_{k-1}\delta_{k-1}\gamma_{k-1}+\lambda^{c}(1-\theta_{k-1})\delta_{k-1}$, $\bar{\bar{\theta}}_{k-1}=(1-\theta_{k-1})(\lambda^{c}\delta_{k-1}-{\lambda^{c}}^{2}\delta_{k-1})$, $\varsigma_{k-1}=\operatorname{\mathsf{E}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$, and $\Xi_{k-1}=\operatorname{\mathsf{cov}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]+2\operatorname{\mathsf{cov}}[\mu_{k-1},(1-\delta_{k-1}\gamma_{k-1})A_{k-1}\tilde{e}_{k-1}|\mathcal{I}^{e}_{k-1},\delta_{k-1},\theta_{k-1}]$.
@@
Observe that we can obtain (29) by plugging (21) into (20), and then subtracting (3) from the result. Suppose for now that $\mu_{k-1}=0$. When feedback information is successfully delivered in the backward channel at time $k$, we have $z^{b}_{k}=\gamma_{k-1}$. Taking the conditional expectation and the conditional covariance of (29) given $\mathcal{I}^{e}_{k}$, when $z^{b}_{k}=\gamma_{k-1}$, i.e., when $\delta_{k-1}=1\wedge\theta_{k-1}=1$, we get
$\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k}\Big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\sigma_{k-1})A_{k-1}\operatorname{\mathsf{E}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}+K_{k}\nu_{k}$ (32)
$\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k}\Big{|}\mathcal{I}^{e}_{k}\Big{]}$ $=(1-\sigma_{k-1})A_{k-1}\operatorname{\mathsf{cov}}\Big{[}\tilde{e}_{k-1}\big{|}\mathcal{I}^{e}_{k-1}\Big{]}A_{k-1}^{T}$ (33)

### v2

@@ VI-D Scalar Sources
Figure 2: Causal tradeoff curves for remote state estimation when $\lambda=0.20$ and $\rho\in\{0.01,0.30,0.70,1.00\}$.
@@
Figure 3: Causal tradeoff curves for remote state estimation when $\lambda=0.40$ and $\rho\in\{0.01,0.30,0.70,1.00\}$.
Figure 4: Causal tradeoff curves for remote state estimation when $\lambda=0.60$ and $\rho\in\{0.01,0.30,0.70,1.00\}$.
Figure 5: Causal tradeoff curves for remote state estimation when $\lambda=0.80$ and $\rho\in\{0.01,0.30,0.70,1.00\}$.
Figure 6: Key trajectories in remote state estimation over a highly unreliable channel without feedback.
Figure 7: Key trajectories in remote state estimation over a highly unreliable channel with unreliable feedback.
Under this setting, the optimal policy profile and the complexity of its computation are provided by the following proposition and corollary.
Proposition 4
The causal tradeoff between the packet rate and the mean square error for a scalar source under unreliable communication with unreliable acknowledgements admits a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ such that $\epsilon^{\star}$ is a symmetric threshold-based scheduling policy given by
$u_{k}=\mathds{1}_{\chi_{k}-\alpha_{k}\geq 0},$ (50)
for $k\in\mathbb{N}_{[0,N-1]}$, in conjunction with $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$, where $\chi_{k}(\breve{e}_{k},R_{k})=\lambda^{c}A_{k}^{2}\breve{e}_{k}^{2}+\lambda^{c}A_{k}^{2}R_{k}+\Delta_{k}$ is a symmetric function of $\breve{e}_{k}$ and also a function of $R_{k}$, which requires solving
$\check{x}_{k}$ $=A_{k-1}\check{x}_{k-1}+\frac{Q_{k}C_{k}}{V_{k}}\nu_{k},$ (51)
$Q_{k}$ $=\frac{1}{\frac{1}{A_{k-1}^{2}Q_{k-1}+W_{k-1}}+\frac{C_{k}^{2}}{V_{k}}}\>,$ (52)
$\breve{e}_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}\breve{e}_{k-1}+\frac{Q_{k}C_{k}}{V_{k}}\nu_{k},$ (53)
$R_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}^{2}R_{k-1}+\bar{\bar{\theta}}_{k-1}A_{k-1}^{2}\breve{e}_{k-1}^{2},$ (54)
for $k\in\mathbb{N}_{[1,N]}$ with initial conditions $\check{x}_{0}=m_{0}+K_{0}\nu_{0}$, $Q_{0}=1/(1/M_{0}+C_{0}^{2}/V_{0})$, $\breve{e}_{0}=Q_{0}C_{0}\nu_{0}/V_{0}$, and $R_{0}=0$, where $\nu_{k}=y_{k}-C_{k}A_{k-1}\check{x}_{k-1}$, $\bar{\theta}_{k-1}=\theta_{k-1}u_{k-1}\gamma_{k-1}+\lambda^{c}(1-\theta_{k-1})u_{k-1}$, and $\bar{\bar{\theta}}_{k-1}=(\lambda^{c}-{\lambda^{c}}^{2})(1-\theta_{k-1})u_{k-1}$; and $\delta^{\star}$ is a linear switching estimation policy given by
$\hat{x}_{k}=A_{k-1}\hat{x}_{k-1}+u_{k-1}\gamma_{k-1}A_{k-1}\tilde{e}_{k-1},$ (55)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}$, where $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, without being influenced by implicit information.
Proof:
These results follow directly from Theorem 1 by writing all equations in a scalar form. ∎
Corollary 5
Let $\breve{e}_{k}$ and $R_{k}$ be discretised in grids with $d_{1}$ and $d_{2}$ points, respectively, and $\chi_{k}$ be evaluated based on Monte Carlo integration with $d_{0}$ samples. The complexity of computing the policy profile $(\epsilon^{\star},\delta^{\star})$ in Proposition 4 is $\mathcal{O}(Nd_{0}d_{1}d_{2})$.
Proof:
[…]
Remark 11
In the scalar regime, matrix-valued quantities reduce to scalars, and inner products become simple multiplications. This leads to a substantial simplification of the recursive equations and the computation of intermediate terms. Moreover, the computational complexity of the optimal policy profile decreases significantly, as the discretisation now only spans one-dimensional spaces. Consequently, scalar systems can serve as a testbed for exploring advanced design strategies before extending them to higher-dimensional settings. In spite of their simplicity, scalar systems carry the key elements of the general theory, making them valuable for theoretical validation and practical prototyping.
VII Numerical Analysis
In this section, we present two illustrative examples that demonstrate the applicability of the theoretical results developed in the paper. The first example involves an industrial machine whose temperature should be monitored at a data centre, while the second example involves a spin-stabilised spacecraft whose angular velocity should be tracked at a ground station. These examples are chosen to reflect practical remote estimation scenarios where communication constraints play a critical role.
VII-A Remotely Tracking Temperature of a Machine
In this example, we consider an industrial machine whose temperature deviation from a known reference temperature evolves according to
$\frac{dT}{dt}=-cT+\zeta,$ (56)
where $T$ is the temperature deviation, $c$ is the heat coefficient, and $\zeta$ is a Gaussian noise. We assume that $c=1.054$. A sensor deployed in the machine measures its temperature. The sensory information should be transmitted over a costly communication channel to a data centre where the temperature is estimated in real-time. The dynamic equation in (56) can be discretised over the time horizon of the experiment $N=1000$. The discrete-time state and output equations of the form (1) and (2) are specified by $A_{k}=0.9$, $C_{k}=1$, $W_{k}=3$, $V_{k}=1$, for $k\in\mathbb{N}_{[0,N]}$, $m_{0}=0$, and $M_{0}=1$. The forward error rate $\lambda$, the backward error rate $\rho$, and the weighting coefficient $\alpha_{k}$ are assumed to be constant over time. We used Monte Carlo simulation, and analysed how feedback influences the performance limits of the networked system in various communication regimes. Figs. 25 depict the causal tradeoff curves when $\lambda$ is equal to $0.20$, $0.40$, $0.60$, and $0.80$, respectively, while $\rho$ takes values in the set $\{0.01,0.30,0.70,1.00\}$. These tradeoff curves delineate the boundaries of achievable regions: any point above a given curve corresponds to a feasible combination of communication cost and estimation error, whereas points below the curve are unattainable under the given condition. From these plots, several key insights emerge. First, for each fixed value of $\lambda$, decreasing $\rho$ (i.e., improving feedback quality) leads to a larger achievable region, enabling better estimation performance with similar communication effort. Second, when $\rho=0.01$ (i.e., feedback is nearly perfect), the encoder has an accurate knowledge of the decoder’s belief, which enhances scheduling decisions and leads to the best system performance. Third, as $\rho$ increases toward $1$ (i.e., feedback becomes unavailable), the encoder receives little useful feedback, resulting in degraded performance and greater packet rates. Lastly, as the forward error rate $\lambda$ increases, the estimation performance degrades and curves shift to the right for all values of $\rho$, illustrating the critical role of reliable forward communication in maintaining estimation quality. These observations underscore the importance of robust communication mechanisms in remote estimation systems and highlight how the optimal policies can adaptively balance communication and estimation objectives even under significant uncertainty.
VII-B Remotely Tracking Angular Velocity of a Spacecraft
In this example, we consider a spin-stabilised spacecraft whose body is spinning about the axis of symmetry, with a constant angular velocity $\omega_{z}=\omega_{0}$. For such a spacecraft, the Euler equation is written as
$\begin{bmatrix}\frac{d\omega_{x}}{dt}\\ \frac{d\omega_{y}}{dt}\\ \frac{d\omega_{z}}{dt}\end{bmatrix}=\begin{bmatrix}0&\frac{I_{y}-I_{z}}{I_{x}}\omega_{0}&0\\ \frac{I_{z}-I_{x}}{I_{y}}\omega_{0}&0&0\\ 0&0&0\end{bmatrix}\begin{bmatrix}\omega_{x}\\ \omega_{y}\\ \omega_{z}\end{bmatrix}+\begin{bmatrix}e_{x}\\ e_{y}\\ e_{z}\end{bmatrix}$, (57)
where $(\omega_{x},\omega_{y},\omega_{z})$ is the angular velocity, $(I_{x},I_{y},I_{z})$ is the moment of inertia, and $(e_{x},e_{y},e_{z})$ is a Gaussian disturbance torque acting on the spacecraft. Note that for spin stability, the spacecraft must be spinning either about the major or minor axis of inertia. We assume that $I_{x}=I_{y}=20\ \text{kg.m${}^{2}$}$, $I_{z}=200\ \text{kg.m${}^{2}$}$, and $\omega_{0}=4\pi\ \text{rad/s}$. A sensor in the spacecraft measures each component of the angular velocity. The sensory information should be transmitted over a costly communication channel to a ground station where the angular velocity is estimated in real-time. The Euler equation in (57) can be discretised over the time horizon of the experiment $N=1000$. The discrete-time state and output equations of the form (1) and (2) are specified by $A_{k}=[0.4258,-0.9048,0;0.9048,0.4258,0;0,0,1]$, $W_{k}=10^{-6}\operatorname{diag}\{0.2245,0.2245,0.0025\}$, $C_{k}=\operatorname{diag}\{1,1,1\}$, $V_{k}=10^{-3}\operatorname{diag}\{1,1,1\}$, for all $k\in\mathbb{N}_{[0,N]}$, $m_{0}=[0;0;2\pi]$, and $M_{0}=10W$. We assume that the weighting coefficient is $\alpha_{k}=6\times 10^{-7}$ for $k\in\mathbb{N}_{[0,N]}$. We studied two communication regimes, and analysed key trajectories of the networked system. The first regime is specified by $\lambda=0.4$ and $\rho=1.0$ (i.e., a highly unreliable channel without feedback), while the second regime by $\lambda=0.4$ and $\rho=0.4$ (i.e., a highly unreliable channel with unreliable feedback). For a simulated realisation of the system, Fig. 6 shows the mean square error, packet transmission, and forward packet loss trajectories under the first regime. In this case, the total mean square error is $0.0486$, with $443$ total transmissions and $231$ forward packet losses. For the same realisation of the system, Fig. 7 shows the mean square error, packet transmission, forward packet loss, and backward packet loss trajectories under the second regime. In this case, the total mean square error is slightly lower at $0.0479$, with $377$ total transmissions, $188$ forward packet losses, and $189$ backward packet losses. As expected, the lack of a feedback channel impairs the encoder’s knowledge of the decoder’s belief, leading to more frequent packet transmissions. However, when feedback is available, even unreliably, the number of packet transmissions decreases significantly while maintaining comparable estimation performance. This demonstrates the critical role of feedback in reducing communication cost in remote estimation. For spacecraft operating in deep-space missions, where communication resources are limited and packet loss is common, these findings suggest that incorporating even partial feedback can be substantially beneficial.
VIII Conclusions
In this article, we developed a comprehensive theoretical framework for remote estimation over unreliable channels with unreliable feedback. We formulated a causal tradeoff between the packet rate and the mean square error, and characterised a globally optimal coding policy profile. We proved that the value function can be transformed into a lower-dimensional form. This dimensionality reduction not only enhances interpretability but also leads to tractable implementation. We proposed an algorithm for the computation of the optimal policies, along with an analysis of its computational complexity. To illustrate the scope of the theoretical results, we examined several special cases that emerge as direct consequences of the general framework. In particular, we studied: reliable channels without feedback, unreliable channels with reliable feedback, unreliable channels without feedback, and scalar sources. We also presented a numerical analysis, demonstrating how the proposed policies perform under different regimes. Our numerical results highlighted the impact of channel reliability on remote estimation, underscoring the critical role of feedback in achieving efficient system operation.
We acknowledge that grid-based discretisation inevitably suffers from the curse of dimensionality as the system dimension increases. For high-dimensional systems, exact implementation may therefore become computationally prohibitive. An important direction for future work is the development of scalable approximation methods, such as function-approximation-based value iteration, adaptive gridding, or learning-based approaches, to enable efficient policy synthesis in large-scale settings while preserving the structural insights established in this paper. We further acknowledge that ignoring quantisation error effectively separates rate-distortion considerations from the estimation problem. Relaxing this assumption would necessitate a joint treatment of quantisation, scheduling, and estimation, which may result in significantly greater analytical and computational complexity. Exploring such integrated designs, along with developing tractable approximation methods for quantised communication, constitutes a natural avenue for future research. Several other promising directions remain open. For instance, analysing the impact of transmission delays in both forward and backward channels could provide a more realistic extension of the current framework. Another important topic is the study of multi-user systems where multiple encoders share a common unreliable channel with unreliable feedback, introducing new challenges in scheduling, resource allocation, and coordination among users.
@@ VIII Conclusions
This section is dedicated to the derivation of the main results. We begin with the following Lemma, showing that the conditional means $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ and $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$ are instrumental in the operations performed by the encoder and the decoder.
@@
Let $\epsilon^{o}$ be a globally optimal scheduling policy that is implemented. Given all information available at the decoder, $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$ is the optimal value that minimises the mean square error at the decoder at time $k$. Moreover, $\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ fuses all the previous and current outputs of the process that are available at the encoder at time $k$. This implies that, from the minimum mean-square-error perspective, the decoder is able to develop a state estimate upon the receipt of a message at time $k+1$, that would be the same if it had all the previous outputs of the process until time $k$, which is the best possible case for the decoder. Hence, both claims hold. ∎
The next two lemmas characterise the optimal estimators at the encoder and the decoder.
@@
The optimal estimator minimising the mean square error at the encoder satisfies the recursive equations
$\check{x}_{k}$ $=m_{k}+K_{k}(y_{k}-C_{k}m_{k}),$ (58)
$m_{k}$ $=A_{k-1}\check{x}_{k-1},$ (59)
$Q_{k}$ $=(M_{k}^{-1}+C_{k}^{T}V_{k}^{-1}C_{k})^{-1},$ (60)
$M_{k}$ $=A_{k-1}Q_{k-1}A_{k-1}^{T}+W_{k-1},$ (61)
@@
The optimal estimator minimising the mean square error at the decoder satisfies the recursive equation
$\hat{x}_{k}$ $=A_{k-1}\hat{x}_{k-1}+u_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1})$
$\quad+(1-u_{k-1})\imath_{k-1}+u_{k-1}(1-\gamma_{k-1})\jmath_{k-1},$ (62)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}$, where $\imath_{k-1}=A_{k-1}\operatorname{\mathsf{E}}[\hat{e}_{k-1}|\mathcal{I}_{k-1}^{d},u_{k-1}=0]$ and $\jmath_{k-1}=A_{k-1}\operatorname{\mathsf{E}}[\hat{e}_{k-1}|\mathcal{I}_{k-1}^{d},u_{k-1}=1]$.
@@
$\operatorname{\mathsf{E}}\Big[x_{k}\big|\mathcal{I}^{d}_{k}\Big]=A_{k-1}\operatorname{\mathsf{E}}\Big[x_{k-1}\big|\mathcal{I}_{k}^{d}\Big],$ (63)
for $k\in\mathbb{N}_{[1,N]}$ as $w_{k-1}$ is independent of $\mathcal{I}^{d}_{k}$ and has zero mean. When sensory information is successfully delivered in the forward channel at time $k$, we have $z^{f}_{k}=\check{x}_{k-1}$. In this case, we get $\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k}]=\operatorname{\mathsf{E}}[x_{k-1}|\check{x}_{k-1},Q_{k-1}]=\check{x}_{k-1}$ as $\{\check{x}_{k-1},Q_{k-1}\}$ is a sufficient statistic of $\mathcal{I}^{d}_{k}$. Hence, using (63), when $z^{f}_{k}=\check{x}_{k-1}$, i.e., $u_{k-1}=1\wedge\gamma_{k-1}=1$, we get
$\operatorname{\mathsf{E}}\Big[x_{k}\big|\mathcal{I}^{d}_{k}\Big]=A_{k-1}\check{x}_{k-1},$ (64)
for $k\in\mathbb{N}_{[1,N]}$. However, when no sensory information is delivered in the forward channel at time $k$, we have $z^{f}_{k}=\mathfrak{E}$ or $z^{f}_{k}=\mathfrak{F}$. Define $p_{k-1}:=\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1},u_{k-1}=0]-\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1}]$ when $z^{f}_{k}=\mathfrak{E}$, and $q_{k-1}:=\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1},u_{k-1}=1]-\operatorname{\mathsf{E}}[x_{k-1}|\mathcal{I}^{d}_{k-1}]$ when $z^{f}_{k}=\mathfrak{F}$. Then, using (63) and the definition of $p_{k-1}$, when $z^{f}_{k}=\mathfrak{E}$, i.e., $u_{k-1}=0$, we get
$\operatorname{\mathsf{E}}\Big[x_{k}\big|\mathcal{I}^{d}_{k}\Big]=A_{k-1}\operatorname{\mathsf{E}}\Big[x_{k-1}\big|\mathcal{I}^{d}_{k-1}\Big]+A_{k-1}p_{k-1},$ (65)
for $k\in\mathbb{N}_{[1,N]}$, where we used the fact that $\{\mathcal{I}^{d}_{k-1},z^{f}_{k}=\mathfrak{E}\}$ is a sufficient statistic of $\mathcal{I}^{d}_{k}$. Moreover, using (63) and the definition of $q_{k-1}$, when $z^{f}_{k}=\mathfrak{F}$, i.e., $u_{k-1}=1\wedge\gamma_{k-1}=0$, we get
$\operatorname{\mathsf{E}}\Big[x_{k}\big|\mathcal{I}^{d}_{k}\Big]=A_{k-1}\operatorname{\mathsf{E}}\Big[x_{k-1}\big|\mathcal{I}^{d}_{k-1}\Big]+A_{k-1}q_{k-1},$ (66)
for $k\in\mathbb{N}_{[1,N]}$, where we used the facts that $\{\mathcal{I}^{d}_{k-1},z^{f}_{k}=\mathfrak{F}\}$ is a sufficient statistic of $\mathcal{I}^{d}_{k}$ and that $x_{k-1}$ is independent of $\gamma_{k-1}$. Now, define $\imath_{k-1}:=A_{k-1}p_{k-1}$ and $\jmath_{k-1}:=A_{k-1}q_{k-1}$ as the signalling residuals. We can obtain (3) by combining (64), (65), and (66). Finally, the initial condition is $\operatorname{\mathsf{E}}[x_{0}]=m_{0}$, as no information is available at time $k=0$. ∎
The next two lemmas provide certain properties pertaining to the estimation mismatch and the signalling residuals $\imath_{k}$ and $\jmath_{k}$.
@@
$\tilde{e}_{k}$ $=(1-u_{k-1}\gamma_{k-1})A_{k-1}\tilde{e}_{k-1}+K_{k}\nu_{k}+\mu_{k-1},$ (67)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\tilde{e}_{0}=K_{0}\nu_{0}$, where $\mu_{k-1}=-(1-u_{k-1})\imath_{k-1}-u_{k-1}(1-\gamma_{k-1})\jmath_{k-1}$. Besides, the conditional mean and the conditional covariance of the estimation mismatch at the encoder satisfy the recursive equations
$\breve{e}_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}\breve{e}_{k-1}+K_{k}\nu_{k}+\varsigma_{k-1},$ (68)
@@
$\quad+\bar{\bar{\theta}}_{k-1}A_{k-1}\breve{e}_{k-1}\breve{e}_{k-1}^{T}A_{k-1}^{T}+\Xi_{k-1},$ (69)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\breve{e}_{0}=K_{0}\nu_{0}$ and $R_{0}=0$, where $\bar{\theta}_{k-1}=\theta_{k-1}u_{k-1}\gamma_{k-1}+\lambda^{c}(1-\theta_{k-1})u_{k-1}$, $\bar{\bar{\theta}}_{k-1}=(\lambda^{c}-{\lambda^{c}}^{2})(1-\theta_{k-1})u_{k-1}$, $\varsigma_{k-1}=\operatorname{\mathsf{E}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},u_{k-1},u_{k-1}\gamma_{k-1}=0]$, and $\Xi_{k-1}=\operatorname{\mathsf{cov}}[\mu_{k-1}|\mathcal{I}^{e}_{k-1},u_{k-1},u_{k-1}\gamma_{k-1}=0]+2\operatorname{\mathsf{cov}}[\mu_{k-1},A_{k-1}\tilde{e}_{k-1}|\mathcal{I}^{e}_{k-1},u_{k-1},u_{k-1}\gamma_{k-1}=0]$.
@@
Observe that we can obtain (67) by plugging (59) into (58), and then subtracting (3) from the result.
@@
Suppose for now that $\mu_{k-1}=0$. When feedback information is successfully delivered in the backward channel at time $k$, we have $z^{b}_{k}=\gamma_{k-1}$. Taking the conditional expectation and the conditional covariance of (67) given $\mathcal{I}^{e}_{k}$, when $z^{b}_{k}=\gamma_{k-1}$, i.e., $u_{k-1}=1\wedge\theta_{k-1}=1$, we get
$\operatorname{\mathsf{E}}\Big[\tilde{e}_{k}\Big|\mathcal{I}^{e}_{k}\Big]$ $=(1\!-\!\sigma_{k-1})A_{k-1}\operatorname{\mathsf{E}}\Big[\tilde{e}_{k-1}\big|\mathcal{I}^{e}_{k-1}\Big]\!+\!K_{k}\nu_{k},$ (70)
$\operatorname{\mathsf{cov}}\Big[\tilde{e}_{k}\Big|\mathcal{I}^{e}_{k}\Big]$ $=(1\!-\!\sigma_{k-1})A_{k-1}\operatorname{\mathsf{cov}}\Big[\tilde{e}_{k-1}\big|\mathcal{I}^{e}_{k-1}\Big]A_{k-1}^{T},$ (71)